OpenStudy (anonymous):

multiply 2(sqrt)11 (times) 8(sqrt) 22

OpenStudy (anonymous):

What have you got so far? \[a\sqrt{b} \times c\sqrt{d}= a*c\sqrt{b*d}\]

jhonyy9 (jhonyy9):

16sqrt(11*22)

OpenStudy (anonymous):

okay i no you suppose to multiply the outside with the out and the inside with the inside okay when i did that i got 14(sqrt)242 and i think thats wrong

OpenStudy (anonymous):

wait 16

OpenStudy (anonymous):

16(sqrt)242 is right. now we need to simplify the sqrt what are factors of 242?

OpenStudy (anonymous):

11 (times)22

OpenStudy (anonymous):

yes what else? 11 and 22 can't be written as a^2

OpenStudy (anonymous):

what can we write 22 as instead?

OpenStudy (anonymous):

2 (times) 11

OpenStudy (anonymous):

good so 242=2*11*11 now what will your answer be?

OpenStudy (anonymous):

i wanna say\[8\sqrt{11}\]

OpenStudy (anonymous):

okay so we have \[16\sqrt{242}=16\sqrt{11}\sqrt{11}\sqrt{2}\]\[=16*11\sqrt{2}\]\[=176\sqrt{2}\]

OpenStudy (anonymous):

\[\sqrt{a}*\sqrt{a}=a\]

OpenStudy (anonymous):

ohh okay 16 (times)11 and what every is in the (sqrt) you just combine it

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

okay thank you ima try the next problem and see what i get